Show that for two scalar fields f and g:
∇.[∇f ×( f ∇g)]= 0
1
Expert's answer
2019-04-30T11:40:26-0400

This is most easily proved by using the Kronecker totally anti-symmetric symbol "\epsilon_{ijk}" such that "\epsilon_{123} = 1". Given two vectors "a" and "b" with the Cartesian components, respectively, "a_i" and "b_i", the Cartesian components of their vector product are given by


"\left[ \bm a \times \bm b \right]_i = \sum_{ij} \epsilon_{ijk} a_j b_k"

and their scalar product is "a \cdot b = \sum_i a_i b_i". Applying these formulas, and using the simple relation "f {\nabla} f = \frac12 {\nabla} f^2", we have


"{\nabla} \cdot \left [ {\nabla} f \times \left ( f {\nabla} g \right ) \right ] = {\nabla} \cdot \left [ f {\nabla} f \times { \nabla} g \right ] = \frac{1}{2} { \nabla} \cdot \left [{ \nabla} f^2 \times { \nabla} g \right ] =""\frac{1}{2} \sum_i \partial_i \left [ {\nabla} f^2 \times { \nabla} g \right ]_i = \frac{1}{2} \sum_{ijk} \partial_i \left ( \epsilon_{ijk} \partial_j f^2 \partial_k g \right ) =""\frac{1}{2} \sum_{ ijk } \epsilon_{ ijk } \left( \partial_i \partial_j f^2 \partial_k g + \partial_j f^2 \partial_i \partial_k g \right ) = 0\, ."

The last equality is valid because is symmetric with respect to {ij}, and is symmetric with respect to {ik}, whereas is totally anti-symmetric.